I could really use some help with this problem:
3x^2 + 6 3/4 x + 6 3/4 = 1 1/2 x^2 + 2 1/4 x
Any help would be greatly appreciated!!!

Answer by stanbon(75887)

3x^2 + 6 3/4 x + 6 3/4 = 1 1/2 x^2 + 2 1/4 x
-----------
3x^2 + (27/4)x + (27/4) = (3/2)x^2 + (9/4)x
-----------
Multiply thru by 4 to get rid of the fractions:
12x^2 + 27x + 27 = 6x^2 + 9x
------
Combine:
6x^2 + 18x + 27 = 0
----
2x^2 + 6x + 9 = 0
----
x = [-6 +- sqrt(36 - 4*2*9)]/(4)
---
x = [-6 +- sqrt(-36)]/4
---
x = [-6 +- 6i]/4
---
x = (-3/2) +- (3/2)i
=========================
Cheers,
Stan H.
============